### v3

@@ D.2 The Case of $\mu_{h}>0$
$\Gamma_{t\in\left\{0\right\}\cup\left[T\right]}=\prod_{s=1}^{t}(1+\mu_{h}\eta_{s})=\begin{cases}\left(1+\frac{1}{\eta+2\kappa_{h}}\right)^{t}&t\leq\tau\\ \left(1+\frac{1}{\eta+2\kappa_{h}}\right)^{\tau}\frac{(t-\tau+1+4\kappa_{h})(t-\tau+2+4\kappa_{h})}{(1+4\kappa_{h})(2+4\kappa_{h})}&t\geq\tau+1\end{cases}.$
So we can obtain
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]$
$\leq$ $\frac{\mu_{h}D_{\psi}(x,x^{1})}{\Gamma_{T}-1}+2\mu_{h}(M^{2}+\sigma^{2})\sum_{t=1}^{T}\frac{\eta_{t}^{2}}{\Gamma_{T}/\Gamma_{t-1}-1}$
$=$ $\frac{\mu_{h}D_{\psi}(x,x^{1})}{\left(1+\frac{1}{\eta+2\kappa_{h}}\right)^{\tau}\frac{(T-\tau+1+4\kappa_{h})(T-\tau+2+4\kappa_{h})}{(1+4\kappa_{h})(2+4\kappa_{h})}-1}+2\mu_{h}(M^{2}+\sigma^{2})\sum_{t=1}^{T}\frac{\eta_{t}^{2}}{\Gamma_{T}/\Gamma_{t-1}-1}$
$\overset{(a)}{\leq}$ $\frac{\mu_{h}D_{\psi}(x,x^{1})}{\left(1+\frac{1}{\eta+2\kappa_{h}}\right)^{T/2}-1}+2\mu_{h}(M^{2}+\sigma^{2})\sum_{t=1}^{T}\frac{\eta_{t}^{2}}{\Gamma_{T}/\Gamma_{t-1}-1}$
$\overset{(b)}{\leq}$ $\frac{\mu_{h}D_{\psi}(x,x^{1})}{\exp\left(\frac{T}{2(1+\eta+2\kappa_{h})}\right)-1}+2\mu_{h}(M^{2}+\sigma^{2})\sum_{t=1}^{T}\frac{\eta_{t}^{2}}{\Gamma_{T}/\Gamma_{t-1}-1}$
$=$ $\frac{\mu_{h}D_{\psi}(x,x^{1})}{\exp\left(\frac{T}{2(1+\eta+2\kappa_{h})}\right)-1}+2\mu_{h}(M^{2}+\sigma^{2})\left(\underbrace{\sum_{t=1}^{\tau}\frac{\eta_{t}^{2}}{\Gamma_{T}/\Gamma_{t-1}-1}}_{\text{I}}+\underbrace{\sum_{t=\tau+1}^{T}\frac{\eta_{t}^{2}}{\Gamma_{T}/\Gamma_{t-1}-1}}_{\text{II}}\right),$ (51)
where $(a)$ is by $T-\tau\geq 0$ and $\tau\geq\frac{T}{2}$, $(b)$ is due to
$\left(1+\frac{1}{\eta+2\kappa_{h}}\right)^{T/2}=\exp\left(\frac{T}{2}\log\left(1+\frac{1}{\eta+2\kappa_{h}}\right)\right)\geq\exp\left(\frac{T}{2(1+\eta+2\kappa_{h})}\right).$
Now we bound
I $=\frac{1}{\mu_{h}^{2}(\eta+2\kappa_{h})^{2}}\sum_{t=1}^{\tau}\frac{1}{\left(1+\frac{1}{\eta+2\kappa_{h}}\right)^{\tau-t+1}\frac{(T-\tau+1+4\kappa_{h})(T-\tau+2+4\kappa_{h})}{(1+4\kappa_{h})(2+4\kappa_{h})}-1}$
$\overset{(c)}{\leq}\frac{1}{\mu_{h}^{2}(\eta+2\kappa_{h})^{2}\left(1+\frac{1}{\eta+2\kappa_{h}}\right)}\sum_{t=1}^{\tau}\frac{1}{\frac{(T-\tau+1+4\kappa_{h})(T-\tau+2+4\kappa_{h})}{(1+4\kappa_{h})(2+4\kappa_{h})}-1}$
$=\frac{1}{\mu_{h}^{2}(\eta+2\kappa_{h})(\eta+1+2\kappa_{h})}\sum_{t=1}^{\tau}\frac{(1+4\kappa_{h})(2+4\kappa_{h})}{(T-\tau)(T-\tau+8\kappa_{h}+3)}$
$=\frac{(1+4\kappa_{h})(2+4\kappa_{h})}{\mu_{h}^{2}(\eta+2\kappa_{h})(\eta+1+2\kappa_{h})}\cdot\frac{\tau}{(T-\tau)(T-\tau+8\kappa_{h}+3)}$
$\overset{(d)}{\leq}\frac{2(1+4\kappa_{h})}{\mu_{h}^{2}(\eta+2\kappa_{h})}\cdot\frac{2(T+1)}{(T-1)(T+5+16\kappa_{h})}$
$\overset{(e)}{\leq}\frac{2\left(2+\frac{1}{\eta+2\kappa_{h}}\right)}{\mu_{h}^{2}}\cdot\frac{6}{T+5+16\kappa_{h})}=\frac{12\left(2+\frac{1}{\eta+2\kappa_{h}}\right)}{\mu_{h}^{2}(T+5+16\kappa_{h})},$ (52)
where $(c)$ is by $\tau-t\geq 0$ and $1+\frac{1}{\eta+2\kappa_{h}}\geq 1$. We use $\eta\geq 0$, $\tau\leq\frac{T+1}{2}$ in $(d)$ and $T\geq 2$ in $(e)$. Next, there is
II $=\frac{4}{\mu_{h}^{2}}\sum_{t=\tau+1}^{T}\frac{1}{(t-\tau+4\kappa_{h})^{2}\left(\frac{(T-\tau+1+4\kappa_{h})(T-\tau+2+4\kappa_{h})}{(t-\tau+4\kappa_{h})(t-\tau+1+4\kappa_{h})}-1\right)}$
$=\frac{4}{\mu_{h}^{2}}\sum_{t=1}^{T-\tau}\frac{1}{(t+4\kappa_{h})^{2}\left(\frac{(T-\tau+1+4\kappa_{h})(T-\tau+2+4\kappa_{h})}{(t+4\kappa_{h})(t+1+4\kappa_{h})}-1\right)}$
$=\frac{4}{\mu_{h}^{2}}\sum_{t=1}^{T-\tau}\frac{t+1+4\kappa_{h}}{(t+4\kappa_{h})\left((T-\tau+1+4\kappa_{h})(T-\tau+2+4\kappa_{h})-(t+4\kappa_{h})(t+1+4\kappa_{h})\right)}$
$\leq\frac{8}{\mu_{h}^{2}}\sum_{t=1}^{T-\tau}\frac{1}{(T-\tau+1+4\kappa_{h})(T-\tau+2+4\kappa_{h})-(t+4\kappa_{h})(t+1+4\kappa_{h})}$
$=\frac{8}{\mu_{h}^{2}}\sum_{t=1}^{T-\tau}\frac{1}{(T-\tau+1-t)(T-\tau+2+8\kappa_{h}+t)}$
$=\frac{8}{\mu_{h}^{2}}\sum_{t=1}^{T-\tau}\frac{1}{2T-2\tau+3+8\kappa_{h}}\left(\frac{1}{T-\tau+1-t}+\frac{1}{T-\tau+2+8\kappa_{h}+t}\right)$
$\leq\frac{8}{\mu_{h}^{2}(2T-2\tau+3+8\kappa_{h})}\left(1+\log(T-\tau)+\log\frac{2T-2\tau+2+8\kappa_{h}}{T-\tau+2+8\kappa_{h}}\right)$
$\leq\frac{8(1+\log T)}{\mu_{h}^{2}(T+2+8\kappa_{h})},$ (53)
where we use $\frac{T}{2}\leq\tau\leq\frac{T+1}{2}$ in the last inequality.
Plugging (52) and (53) into (51) to get
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]$ $\leq\frac{\mu_{h}D_{\psi}(x,x^{1})}{\exp\left(\frac{T}{2(1+\eta+2\kappa_{h})}\right)-1}+\frac{2(M^{2}+\sigma^{2})}{\mu_{h}}\left(\frac{12\left(2+\frac{1}{\eta+2\kappa_{h}}\right)}{T+5+16\kappa_{h}}+\frac{8(1+\log T)}{T+2+8\kappa_{h}}\right)$
$=O\left(\frac{\mu_{h}D_{\psi}(x,x^{1})}{\exp\left(\frac{T}{2(1+\eta+2\kappa_{h})}\right)-1}+\left(1\lor\frac{1}{\eta+2\kappa_{h}}\right)\frac{(M^{2}+\sigma^{2})\log T}{\mu_{h}(T+\kappa_{h})}\right).$ (54)
@@
$F(x^{T+1})-F(x)\leq O\left(\frac{\mu_{h}D_{\psi}(x,x^{1})}{\exp\left(\frac{T}{2(1+\eta+2\kappa_{h})}\right)-1}+\left(1\lor\frac{1}{\eta+2\kappa_{h}}\right)\frac{(M^{2}+\sigma^{2}\log\frac{1}{\delta})\log T}{\mu_{h}(T+\kappa_{h})}\right).$
@@
When $\mu_{f}=0$ and $\mu_{h}>0$, suppose the condition of $\eta_{t\in\left[T\right]}\leq\frac{1}{2L\lor\mu_{f}}=\frac{1}{2L}$ in Lemma 4.3 holds, we will have with probability at least $1-\delta$
@@
Observe that
$\gamma_{t\in\left[T\right]}=\eta_{t}\prod_{s=2}^{t}\frac{1+\mu_{h}\eta_{s-1}}{1-\mu_{f}\eta_{s}}=\eta_{t}\prod_{s=2}^{t}(1+\mu_{h}\eta_{s-1})=\eta_{t}\Gamma_{t-1}=\frac{\Gamma_{t}-\Gamma_{t-1}}{\mu_{h}},$
where $\Gamma_{t\in\left\{0\right\}\cup\left[T\right]}\coloneqq\prod_{s=1}^{t}(1+\mu_{h}\eta_{s})$. So (55) can be rewritten as
$F(x^{T+1})-F(x)\leq\frac{4\mu_{h}D_{\psi}(x,x^{1})}{\Gamma_{T}-1}+4\mu_{h}\left(M^{2}+\sigma^{2}\left(1+2\log\frac{2}{\delta}\right)\right)\sum_{t=1}^{T}\frac{\eta_{t}^{2}}{\Gamma_{T}/\Gamma_{t-1}-1}.$ (56)
We can check that
@@
Therefore, (56) is true for all cases.
@@
$F(x^{T+1})-F(x)\leq O\left(\frac{\mu_{h}D_{\psi}(x,x^{1})}{\exp\left(\frac{T}{2(1+\eta+2\kappa_{h})}\right)-1}+\left(1\lor\frac{1}{\eta+2\kappa_{h}}\right)\frac{(M^{2}+\sigma^{2}\log\frac{1}{\delta})\log T}{\mu_{h}(T+\kappa_{h})}\right).$
@@ Appendix E Unified Theoretical Analysis under Heavy-Tailed Noise
In this section, we present the analysis under heavy-tailed noise. Unlike Lemma 4.1, the requirement of $\eta_{t\in\left[T\right]}\leq\frac{1}{2L}$ is removed.
@@
where $\xi^{t}\coloneqq\widehat{g}^{t}-\mathbb{\mathbb{E}}\left[\widehat{g}^{t}\mid\mathcal{F}^{t-1}\right],\forall t\in\left[T\right]$ and $z^{t}\coloneqq\frac{v_{0}}{v_{t}}x+\sum_{s=1}^{t}\frac{v_{s}-v_{s-1}}{v_{t}}x^{s},\forall t\in\left\{0\right\}\cup\left[T\right]$.
@@
where $g^{t}\coloneqq\mathbb{\mathbb{E}}\left[\widehat{g}^{t}|\mathcal{F}^{t-1}\right]\in\partial f(x^{t})$ and $\xi^{t}=\widehat{g}^{t}-g^{t}$. Next we bound these four terms respectively.
@@
$\sum_{t=1}^{T}\eta_{t}v_{t}\left(F(x^{t+1})-F(z^{t})\right)\geq\eta_{T}v_{T}\left(F(x^{T+1})-F(x)\right),$ (64)
@@
Under Assumptions 2.-4. and 5C. with $\mu_{f}=\mu_{h}=0$, then for any $x\in\mathcal{X}$ we have
$\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq\frac{D_{\psi}(x,x^{1})}{\sum_{t=1}^{T}\eta_{t}}+\frac{(2-p)2^{\frac{3p-4}{2-p}}L^{\frac{p}{2-p}}}{p}\sum_{t=1}^{T}\frac{\eta_{t}^{\frac{p}{2-p}}}{\sum_{s=t}^{T}\eta_{s}}+\frac{4^{p-1}(M^{p}+\sigma^{p})}{p}\sum_{t=1}^{T}\frac{\eta_{t}^{p}}{\sum_{s=t}^{T}\eta_{s}}.$
@@
$\eta_{T}v_{T}\mathbb{\mathbb{E}}\left[F(x^{T+1})-F(x)\right]\leq$ $v_{0}D_{\psi}(x,x^{1})+\sum_{t=1}^{T}\eta_{t}v_{t-1}\mathbb{\mathbb{E}}\left[\left\langle\xi^{t},z^{t-1}-x^{t}\right\rangle\right]$
$+\sum_{t=1}^{T}\frac{\eta_{t}^{\frac{p}{2-p}}v_{t}(2-p)2^{\frac{3p-4}{2-p}}L^{\frac{p}{2-p}}+\eta_{t}^{p}v_{t}4^{p-1}(M^{p}+\mathbb{\mathbb{E}}\left[\left\|\xi^{t}\right\|_{*}^{p}\right])}{p}$
@@
where the last line is due to $\mathbb{\mathbb{E}}\left[\left\|\xi^{t}\right\|_{*}^{p}\right]=\mathbb{\mathbb{E}}\left[\mathbb{\mathbb{E}}\left[\left\|\xi^{t}\right\|_{*}^{p}\mid\mathcal{F}^{t-1}\right]\right]\leq\sigma^{p}$ (Assumption 5C.) and $\mathbb{\mathbb{E}}\left[\left\langle\xi^{t},z^{t-1}-x^{t}\right\rangle\right]=\mathbb{\mathbb{E}}\left[\left\langle\mathbb{\mathbb{E}}\left[\xi^{t}|\mathcal{F}^{t-1}\right],z^{t-1}-x^{t}\right\rangle\right]=0$ ($z^{t-1}-x^{t}\in\mathcal{F}^{t-1}$ and Assumption 4.). Finally, we divide both sides by $\eta_{T}v_{T}$ and plug in $v_{t\in\left\{0\right\}\cup\left[T\right]}=\frac{\eta_{T}}{\sum_{s=t\lor 1}^{T}\eta_{s}}$ to finish the proof. ∎

### v4

@@ D.2 The Case of $\mu_{h}>0$
$\gamma_{t\in\left[T\right]}=\eta_{t}\prod_{s=2}^{t}(1+\mu_{h}\eta_{s-1})=\begin{cases}\eta_{1}(1+\mu_{h}\eta_{1})^{t-1}&t\leq\tau\\ \frac{2(t-\tau+1+4\kappa_{h})}{\mu_{h}(1+4\kappa_{h})(2+4\kappa_{h})}(1+\mu_{h}\eta_{1})^{\tau}&t\geq\tau+1\end{cases}.$ (55)
So, there is
[…]
$\overset{(a)}{\leq}\frac{\mu_{h}}{\exp\left(\frac{\mu_{h}\eta_{1}\tau}{1+\mu_{h}\eta_{1}}\right)-1}=\frac{\mu_{h}}{\exp\left(\frac{\tau}{1+\eta+2\kappa_{h}}\right)-1}\overset{\tau\geq\frac{T}{2}}{\leq}\frac{\mu_{h}}{\exp\left(\frac{T}{2(1+\eta+2\kappa_{h})}\right)-1},$ (56)
where $(a)$ is by
$(1+\mu_{h}\eta_{1})^{\tau}=\exp\left(\tau\log\left(1+\mu_{h}\eta_{1}\right)\right)\geq\exp\left(\frac{\mu_{h}\eta_{1}\tau}{1+\mu_{h}\eta_{1}}\right).$
Now, we observe that
$\sum_{t=1}^{T}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T}\gamma_{s}}=\sum_{t=1}^{\tau}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T}\gamma_{s}}+\sum_{t=\tau+1}^{T}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T}\gamma_{s}}=\underbrace{\sum_{t=1}^{\tau}\frac{\gamma_{t}\eta_{1}}{\sum_{s=t}^{T}\gamma_{s}}}_{\mathrm{I}}+\underbrace{\sum_{t=\tau+1}^{T}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T}\gamma_{s}}}_{\mathrm{II}}.$
We bound terms $\mathrm{I}$ and $\mathrm{II}$ separately in the following.
  • Term $\mathrm{I}$: When $t\leq\tau$, we have
[…]
    $\leq\frac{\eta_{1}}{(1+\mu_{h}\eta_{1})\sum_{s=\tau+1}^{T}\frac{2(s-\tau+1+4\kappa_{h})}{\mu_{h}(1+4\kappa_{h})(2+4\kappa_{h})}}=\frac{\mu_{h}(1+4\kappa_{h})(2+4\kappa_{h})\eta_{1}}{(1+\mu_{h}\eta_{1})(T-\tau)(T-\tau+3+8\kappa_{h})},$
    implying that
    $\mathrm{I}$ $\leq\frac{\mu_{h}(1+4\kappa_{h})(2+4\kappa_{h})\eta_{1}^{2}\tau}{(1+\mu_{h}\eta_{1})(T-\tau)(T-\tau+3+8\kappa_{h})}\overset{\tau\leq\frac{T+1}{2},T\geq 2}{\leq}\frac{6\mu_{h}(1+4\kappa_{h})(2+4\kappa_{h})\eta_{1}^{2}}{(1+\mu_{h}\eta_{1})(T+5+16\kappa_{h})}$
    $=\frac{6(1+4\kappa_{h})(2+4\kappa_{h})}{\mu_{h}(\eta+2\kappa_{h})(1+\eta+2\kappa_{h})(T+5+16\kappa_{h})}\leq\frac{12\left(2+\frac{1}{\eta+2\kappa_{h}}\right)}{\mu_{h}(T+5+16\kappa_{h})}.$
  • Term $\mathrm{II}$: When $t\geq\tau+1$, we have
[…]
    implying that
    $\mathrm{II}$ $\leq\sum_{t=\tau+1}^{T}\frac{8}{\mu_{h}(T-t+1)(T+t-2\tau+2+8\kappa_{h})}$
    $=\frac{8}{\mu_{h}(2T-2\tau+3+8\kappa_{h})}\sum_{t=\tau+1}^{T}\frac{1}{T-t+1}+\frac{1}{T+t-2\tau+2+8\kappa_{h}}$
    $\leq\frac{8}{\mu_{h}(2T-2\tau+3+8\kappa_{h})}\left(1+\log(T-\tau)+\log\frac{2T-2\tau+2+8\kappa_{h}}{T-\tau+2+8\kappa_{h}}\right)$
    $\overset{\frac{T}{2}\leq\tau\leq\frac{T+1}{2}}{\leq}\frac{8(1+\log T)}{\mu_{h}(T+2+8\kappa_{h})}.$
Therefore, we obtain
$\sum_{t=1}^{T}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T}\gamma_{s}}\leq\frac{12\left(2+\frac{1}{\eta+2\kappa_{h}}\right)}{\mu_{h}(T+5+16\kappa_{h})}+\frac{8(1+\log T)}{\mu_{h}(T+2+8\kappa_{h})}\leq\frac{12\left(2+\frac{1}{\eta+2\kappa_{h}}\right)+8(1+\log T)}{\mu_{h}(T+2+8\kappa_{h})}.$ (57)
We combine (52), (56), and (57) to conclude
$\mathbb{E}\left[F(x^{T+1})-F(x)\right]$ $\leq\frac{\mu_{h}D_{\psi}(x,x^{1})}{\exp\left(\frac{T}{2(1+\eta+2\kappa_{h})}\right)-1}+\frac{8(M^{2}+\sigma^{2})\left[3\left(2+\frac{1}{\eta+2\kappa_{h}}\right)+2(1+\log T)\right]}{\mu_{h}(T+2+8\kappa_{h})}$
$=O\left(\frac{\mu_{h}D_{\psi}(x,x^{1})}{\exp\left(\frac{T}{2(1+\eta+2\kappa_{h})}\right)-1}+\left(\frac{1}{\eta+2\kappa_{h}}\lor\log T\right)\frac{(M^{2}+\sigma^{2})}{\mu_{h}(T+\kappa_{h})}\right).$ (58)
@@
$F(x^{T+1})-F(x)\leq O\left(\frac{\mu_{h}D_{\psi}(x,x^{1})}{\exp\left(\frac{T}{2(1+\eta+2\kappa_{h})}\right)-1}+\left(\frac{1}{\eta+2\kappa_{h}}\lor\log T\right)\frac{M^{2}+\sigma^{2}\log\frac{1}{\delta}}{\mu_{h}(T+\kappa_{h})}\right).$
@@
When $\mu_{f}=0$ and $\mu_{h}>0$, suppose the condition of $\eta_{t\in\left[T\right]}\leq\frac{1}{2L\lor\mu_{f}}=\frac{1}{2L}$ in Lemma 4.3 holds, we have, with probability at least $1-\delta$,
@@
where $\gamma_{t\in\left[T\right]}=\eta_{t}\prod_{s=2}^{t}\frac{1+\mu_{h}\eta_{s-1}}{1-\mu_{f}\eta_{s}}=\eta_{t}\prod_{s=2}^{t}(1+\mu_{h}\eta_{s-1})$. Now let us check the condition of $\eta_{t\in\left[T\right]}\leq\frac{1}{2L}$ for the two choices respectively:
@@
Therefore, (59) holds for both cases.
@@
$F(x^{T+1})-F(x)\leq O\left(\frac{\mu_{h}D_{\psi}(x,x^{1})}{\exp\left(\frac{T}{2(1+\eta+2\kappa_{h})}\right)-1}+\left(\frac{1}{\eta+2\kappa_{h}}\lor\log T\right)\frac{M^{2}+\sigma^{2}\log\frac{1}{\delta}}{\mu_{h}(T+\kappa_{h})}\right).$
Remark D.7.
The $O(\log T)$ factor in Theorems D.5 and D.6 could be removed with more sophisticated step sizes, similar to those in Theorems D.3 and D.4. However, we choose not to include them to keep the paper concise.
@@ Appendix E Unified Theoretical Analysis under Heavy-Tailed Noise
In this section, we present the analysis under heavy-tailed noise. We recall that the mirror map $\psi$ is assumed to be $(1,\frac{p}{p-1})$ uniformly convex now, i.e.,
$D_{\psi}(x,y)=\psi(x)-\psi(y)-\left\langle\nabla\psi(y),x-y\right\rangle\geq\frac{p-1}{p}\left\|x-y\right\|^{\frac{p}{p-1}},\forall x,y\in\mathcal{X}.$
We first prove the following general lemma. Unlike Lemma 4.1, the requirement of $\eta_{t\in\left[T\right]}\leq\frac{1}{2L}$ is removed.
@@
where $\xi^{t}\coloneqq\widehat{g}^{t}-\mathbb{E}\left[\widehat{g}^{t}\mid\mathcal{F}^{t-1}\right],\forall t\in\left[T\right]$ and $z^{t}\coloneqq\frac{v_{0}}{v_{t}}x+\sum_{s=1}^{t}\frac{v_{s}-v_{s-1}}{v_{t}}x^{s},\forall t\in\left\{0\right\}\cup\left[T\right]$.
@@
where $g^{t}\coloneqq\mathbb{E}\left[\widehat{g}^{t}|\mathcal{F}^{t-1}\right]\in\partial f(x^{t})$ and $\xi^{t}=\widehat{g}^{t}-g^{t}$. Next we bound these four terms respectively.
@@
$\sum_{t=1}^{T}\eta_{t}v_{t}\left(F(x^{t+1})-F(z^{t})\right)\geq\eta_{T}v_{T}\left(F(x^{T+1})-F(x)\right).$ (67)
@@
Under Assumptions 2.-4. and 5C. with $\mu_{f}=\mu_{h}=0$, then for any $x\in\mathcal{X}$, we have
$\mathbb{E}\left[F(x^{T+1})-F(x)\right]\leq\frac{D_{\psi}(x,x^{1})}{\sum_{t=1}^{T}\eta_{t}}+\frac{(2-p)2^{\frac{3p-4}{2-p}}L^{\frac{p}{2-p}}}{p}\sum_{t=1}^{T}\frac{\eta_{t}^{\frac{p}{2-p}}}{\sum_{s=t}^{T}\eta_{s}}+\frac{4^{p-1}(M^{p}+\sigma^{p})}{p}\sum_{t=1}^{T}\frac{\eta_{t}^{p}}{\sum_{s=t}^{T}\eta_{s}}.$
@@
$\eta_{T}v_{T}\mathbb{E}\left[F(x^{T+1})-F(x)\right]\leq$ $v_{0}D_{\psi}(x,x^{1})+\sum_{t=1}^{T}\eta_{t}v_{t-1}\mathbb{E}\left[\left\langle\xi^{t},z^{t-1}-x^{t}\right\rangle\right]$
$+\sum_{t=1}^{T}\frac{\eta_{t}^{\frac{p}{2-p}}v_{t}(2-p)2^{\frac{3p-4}{2-p}}L^{\frac{p}{2-p}}+\eta_{t}^{p}v_{t}4^{p-1}(M^{p}+\mathbb{E}\left[\left\|\xi^{t}\right\|_{*}^{p}\right])}{p}$
@@
where the last line is due to $\mathbb{E}\left[\left\|\xi^{t}\right\|_{*}^{p}\right]\leq\sigma^{p}$ (Assumption 5C.) and $\mathbb{E}\left[\left\langle\xi^{t},z^{t-1}-x^{t}\right\rangle\right]=\mathbb{E}\left[\left\langle\mathbb{E}\left[\xi^{t}|\mathcal{F}^{t-1}\right],z^{t-1}-x^{t}\right\rangle\right]=0$ ($z^{t-1}-x^{t}\in\mathcal{F}^{t-1}$ and Assumption 4.). Finally, we divide both sides by $\eta_{T}v_{T}$ and plug in $v_{t\in\left\{0\right\}\cup\left[T\right]}=\frac{\eta_{T}}{\sum_{s=t\lor 1}^{T}\eta_{s}}$ to finish the proof. ∎